OpenStudy (anonymous):

What are the factors of x2 + 10x + 24? (x − 12)(x − 2) (x + 12)(x + 2) (x + 4)(x + 6) <- my answer correct me if I'm wrong thxxxx (x − 4)(x − 6)

OpenStudy (anonymous):

there is a simple way to solving this

OpenStudy (anonymous):

1*24=24 4+6=10 4*6=24

OpenStudy (anonymous):

I will check your answer

OpenStudy (anonymous):

you are correct.

OpenStudy (anonymous):

Yes that is the answer....

OpenStudy (anonymous):